An object moves with a velocity of 8.5 m/s [S42W]. It then changes its velocity in the next 2.0 s to 19 m/s [W35N]. What is the acceleration of the object?


1
Expert's answer
2021-02-17T11:01:34-0500
"\bold{a}=\frac{(19\sin{35},19\cos{35})-(8.5\cos{42},-8.5\sin{42})}{2}\\\bold{a}=(2.29,10.63)\frac{m}{s^2}\\a=10.87\frac{m}{s^2}\text{ N78W }"